Please help!
A child goes to the bank to deposit her savings. She has $4 worth of nickels and dimes. If she has 5 more nickels than dimes, how many of each type of coin does she have?
Thank you!

Answer by algebrahouse.com(1659)

x = number of dimes
x + 5 = number of nickels {she has five more nickels than dimes}

10x + 5(x + 5) = 400 {number of coins multiplied by value of coin equals total value}
10x + 5x + 25 = 400 {used distributive property}
15x + 25 = 400 {combined like terms}
15x = 375 {subtracted 25 from both sides}
x = 25 {divided both sides by 15}
x + 5 = 30 {substituted 25, in for x, into x + 5}

25 dimes and 30 nickels

For more help from me, visit: www.algebrahouse.com